OpenStudy (anonymous):

Given: Area of the triangle is 40 and its sides are 40,35,55 cm. find the length of the longest altitude.

OpenStudy (gorv):

1/2*base*height=40

OpenStudy (gorv):

calculate height

OpenStudy (anonymous):

base=?

OpenStudy (anonymous):

\[A=\frac{1}{2}ab\sin\theta\]

OpenStudy (anonymous):

\[\cos\theta=\frac{a^2+b^2-c^2}{2(a)(b)}\]

OpenStudy (gorv):

put all the length in place of base the maximum value of height willbe your answer

OpenStudy (anonymous):

@gorv A=1/2 bh only works for right angled triangles. >.>

OpenStudy (gorv):

@Azteck we have to calculate height and it works for mst of the triangls okkk if i m wrong tell altitude by ur formula??

OpenStudy (anonymous):

But it also applies if you know which side is the base.

OpenStudy (anonymous):

@Azteck A=1/2 bh works for any type of triangle......only if you are given the required lengths......

OpenStudy (gorv):

thats wat i m saying the side which will give max altitude will be considered as base

OpenStudy (gorv):

35 will give max height so take it as base

OpenStudy (anonymous):

how 35 cmes base

OpenStudy (gorv):

put all the sides in formula told base =35 will give max altitude that is height so side=35 will be consideredas the base

OpenStudy (anonymous):

k

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

@gorv thanks bro

OpenStudy (gorv):

take same approch